A photographic "stop bath" contains 160mL of pure acetic acid. HC2H3O2(1) in 650 mL solution. What is the v/v concentration of acetic acid in the stop bath?


1
Expert's answer
2021-10-18T06:49:14-0400

Solution:

acetic acid = HC2H3O2 - solute

Volume of acetic acid = 160 mL

Volume of the solution = 650 mL


v/v concentration = (solute volume / volume of the solution) × 100%

v/v(HC2H3O2) = (160 mL / 650 mL) × 100% = 24.6%

v/v(HC2H3O2) = 24.6%


Answer: The v/v concentration of acetic acid is 24.6%